# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 7.9.2 The Global Rise of Kidney Exchange Programs Since 2008
After the first kidney exchange performed in the U.S. in Rhode Island in 2000, progress in kidney exchange was initially slow. Between 2000 and 2003—prior to our engagement with the transplantation community—only 31 kidney exchange transplants were performed nationwide over four years. In sharp contrast, 4,387 transplants were conducted between 2020 and 2023 (OPTN, 2024a). A record 1,283 kidney exchange transplants were performed in 2023 alone, and by December 2024, the cumulative total exceeded 13,000 patients in the U.S. (For annual transplant numbers from 2000 to 2023, see Figure 7.)
@@
According to estimates by Teltser (2019), 64% of these transplants—approximately 8,300—represent new transplants that would not have occurred without kidney exchange. Furthermore, Teltser (2019) highlights that kidney exchange has also improved average graft survival rates, despite the fact that kidney exchange recipients often come from more disadvantaged groups. These include highly sensitized patients with antibodies against large portions of the population, older patients, and recipients of kidneys from older donors.
@@
Over the past two decades, kidney exchange programs have expanded beyond the U.S., providing life-saving transplants to patients worldwide. Countries such as the U.K., Turkey, Canada, India, Saudi Arabia, the Netherlands, and Australia now perform some of the highest kidney exchange volumes outside the U.S. Altogether, the number of kidney exchange transplants carried out globally to date exceeds 20,000.
@@
The U.K. was the first country outside the U.S. to adopt market design tools for kidney exchange, including the implementation of optimal matching algorithms, and it now operates the largest national system. Since 2007, the National Health Service Blood and Transplant (NHSBT) has coordinated kidney exchanges through the UK Living Donor Kidney Sharing Scheme. This program, which runs every three months, has collaborated with market design researcher David Manlove since 2008 and celebrated its 2000th kidney exchange transplant in September 2024.
@@
Initially limited to 2-way and 3-way cycles (Biró et al., 2009), the U.K. scheme has also been organizing non-directed donor (NDD) chains of up to three recipients since 2012 (Manlove and O’Malley, 2014; Ferrari et al., 2014). Among the 670 kidney exchange transplants conducted between April 2020 and March 2024, 126 (18.8%) were carried out in 2-way cycles, 118 (17.6%) in 3-way cycles, 215 (32.1%) in 2-way NDD chains, and 211 (31.5%) in 3-way NDD chains (NHS Blood and Transplant, 2024).
@@
Two key factors behind the U.K.’s success are the substantial participation of altruistic non-directed donors (NDDs) and compatible donor-recipient pairs. Between April 2015 and March 2024, there were 653 NDDs, with 347 donating directly to the deceased donor list and 306 donating through the kidney sharing scheme (NHS Blood and Transplant, 2024). Over this period, the proportion of NDDs participating in kidney exchanges grew from approximately 30% to 35% to 55% to 65%. Additionally, 119 compatible pairs participated in kidney exchanges during the same period.
@@
Taken together, the U.K.’s centrally managed national system—which incorporates up to 3-way cycles and NDD chains, employs an optimal matching algorithm, and benefits from significant participation by altruistic donors and compatible pairs—stands out as perhaps the closest large-scale realization of the vision pioneered in Roth et al. (2004, 2005b, 2006, 2007).
@@
In Turkey, over 80% of kidney transplants are from living donors, and the country led the world in living kidney donation rates in 2023, exceeding 35 donors per million people (IRODaT, 2024). Performing over 3,000 transplants annually, Turkey also ranks third globally in living-donor kidney transplants (LDKTs), behind only India and the U.S. These factors make Turkey an ideal candidate to implement a national centralized system. Unfortunately, as discussed later in Section 12.4, my efforts with Ünver in this direction have not succeeded so far.
@@
Nevertheless, kidney exchange is widely practiced in Turkey through several single-hospital programs, with more than 2,000 transplants performed as of 2023. However, NDD chains are not utilized, as altruistic donation is prohibited by law.
@@
The Turkish Ministry of Health and the Turkish Society of Nephrology have tracked kidney exchange activity nationwide since 2012 (Turkish Society of Nephrology, 2023). According to their registries, 2,067 kidney exchange transplants were performed between 2012 and 2023, accounting for 6.3% of the country’s total LDKTs. This relatively modest rate reflects the absence of a national kidney exchange system and the limited availability of kidney exchange in only some hospitals. However, in hospitals with kidney exchange programs, rates are significantly higher, around 15%. For example, Altun and Yavuz (2023) documented 141 kidney exchange transplants at Istanbul Göztepe Medicalpark Hospital from July 2011 to June 2021, representing 16.1% of the LDKTs performed at the center during that period.
Kidney exchange activity in Turkey predates 2012, with significant activity reported between 2006 and 2011. According to Yücetin et al. (2013), among the 59 hospitals performing LDKTs nationwide, Antalya Medicalpark Hospital, Istanbul Göztepe Medicalpark Hospital, and Akdeniz University Medical School Transplantation Center reported the highest kidney exchange volumes during this period. Together, they conducted 272 transplants, accounting for 14.4% of the LDKTs carried out at these centers.
Canada’s national Kidney Paired Donation program has been managed by Canadian Blood Services since 2009. The system operates every four months and incorporates cycles and NDD chains of up to 4-way. It is among the most successful kidney exchange programs in the world, largely due to the significant number of NDD chains organized through the contributions of altruistic non-directed donors (Ferrari et al., 2014). In May 2023, it reached its 1,000th kidney exchange transplant, 68% of which were facilitated by donations from 200 altruistic non-directed donors.
@@
India, the most populous country in the world, leads globally in LDKTs, performing over 10,000 transplants in 2022 (IRODaT, 2024). Kidney exchange is widespread and conducted through numerous single-hospital programs across the country. However, as in Turkey, NDD chains are not practiced because altruistic donation is prohibited by law. Unlike in many other countries, kidney exchange operations in India are often carried out over several days, enabling the execution of large exchange cycles.
@@
Kute et al. (2024) reports that the Dr. HL Trivedi Institute of Transplantation Sciences in Ahmedabad—described as the largest single-center kidney exchange program in Asia—performed 539 kidney exchange transplants between January 2000 and March 2024. These exchanges included one 10-way and two 6-way exchanges conducted over multiple days and accounted for 10% of the 5,346 LDKTs performed at the center during this time. Similarly, Mahatma Gandhi Medical College in Jaipur, described by Mangal et al. (2024) as the largest program in Northern India, reported conducting 127 kidney exchange transplants between 2014 and March 2024.
@@
In Saudi Arabia, as in Turkey and many other Asian countries, more than 80% of kidney transplants are from living donors. In 2023, the country ranked second worldwide—after Turkey—in the living kidney donation rate, with more than 35 donors per million people (IRODaT, 2024). The King Faisal Specialist Hospital and Research Centre celebrated its 500th kidney exchange transplant in October 2024 as part of its single-hospital Kidney Paired Donation Program, which has been in operation since 2011.
@@
National kidney exchange programs were launched in Australia and New Zealand in 2010 and 2011, respectively. These efforts were merged in 2019 into the Australian and New Zealand Paired Kidney Exchange (ANZKX) program. Since their inception, these programs have facilitated more than 400 kidney exchange transplants.
In Europe, similar cross-border collaborations have emerged over the past decade, particularly in countries with relatively few living donors. One of the most active initiatives is the Scandiatransplant Kidney Exchange Programme (STEP), which involves transplant centers in Denmark, Finland, Norway, and Sweden (Weinreich et al., 2023).
Launched in 2019 with the advocacy of design economist Tommy Andersson and initially limited to 2-way exchanges, STEP has been organizing up to 3-way exchanges and NDD chains since January 2024. As of September 2024, the program has facilitated a total of 82 transplants (Scandiatransplant, 2024).
Two other cross-border kidney exchange partnerships currently operate in Europe: one between Austria, the Czech Republic, and Israel, and another between Italy, Portugal, and Spain.
Due to the limited size of their national pools, Austria and the Czech Republic—both of which introduced national kidney exchange systems in 2011—merged their programs in 2015. This collaboration led to the first cross-border kidney exchange in 2016 (Böhmig et al., 2017) and was further expanded to include Israeli centers in 2019. Between 2011 and 2019, these systems led to 81 transplants, including 33 after 2015 under the unified program, of which 8 were via cross-border exchanges (Viklicky et al., 2020).
@@
Similarly, Italy, Spain, and Portugal established their national kidney exchange programs in 2006, 2009, and 2010, respectively. In 2017, the three countries formalized a cooperation protocol resulting in their first successful cross-border exchange in 2018 (Valentín et al., 2019). Unlike other multi-country collaborations, however, each country continues to operate its own independent national system before pursuing cross-border matches.
@@
ENCKEP led to a follow-on project called Software for Transnational Kidney Exchange Programmes (KEP-SOFT), funded by a COST Innovators Grant from November 2021 to October 2022. This effort produced KEPsoft, a software solution for national and international kidney exchange (KEP-SOFT, 2024). [^88] KEPsoft was jointly developed by the University of Glasgow, INESC TEC (Porto), HUN-REN-KRTK (Budapest), and Óbuda University (Budapest) (KEP-SOFT, 2024).
Further affirming the growing enthusiasm for kidney exchange in Europe, the European Commission launched a 2023 call for proposals for action grants under its Health Systems and Healthcare Workforce initiative, specifically targeting “Facilitating Organ Paired Exchange” (European Commission, 2023). This initiative aims to harmonize organ donation and transplantation practices across EU member states by developing a shared algorithm for matching donor-patient pairs and establishing common protocols, platforms, and governance frameworks. The ultimate goal is to create an EU-wide system that expands kidney exchange opportunities and improves patient outcomes. [^89] The EURO-KEP project, led by the Spanish National Transplant Organization (ONT), was the sole project funded under this call and is supported by KEPsoft software (KEP-SOFT, 2024).
@@
My holistic approach to research and policy in school choice (see Section 6) and kidney exchange (see Section 7) was primarily driven by instinct. Early in my academic career, I did not consciously follow any specific institutional design framework. In both applications, my goal was to address the limitations of existing resource allocation systems in ways that could gain acceptance from decision makers vested in preserving them.
Clarity began to emerge as I observed parallels between the economist-driven school choice reform in Boston (see Section 6.10) and the natural reforms in England and Chicago (see Section 6.12). It became evident that the methodology evolving from my holistic research and policy efforts could serve as the foundation for a promising institutional design framework on a broader scale.
I interpreted these developments in England and Chicago as providing external validity for this emerging methodology. My approach—uncovering how real-life institutions had diverged from their intended roles and restoring their alignment—resonated with decision makers. This emphasis on aligning institutions with their original intentions enabled me to achieve policy impact, which may explain my success.
@@
In May 2011, I received an unexpected email from (then) Major Tobias Switzer, a U.S. Air Force officer seeking feedback on his master’s thesis (Switzer, 2011) from the Catholic University of Chile. Following the basic structure of Balinski and Sönmez (1999) and adopting the no justified envy axiom (see Definition 5) for his framework, Switzer examined the branching process of cadets into military specialties at the United States Military Academy (USMA). [^90] Following the terminology introduced in Balinski and Sönmez (1999), Switzer (2011) also refers to the no justified envy axiom as fairness.
Upon reading his thesis, I found the problem fascinating and his critique of the mechanism used at West Point—henceforth the USMA-2006 mechanism—to be valid. His thesis marked the start of another “discovery-invention cycle.” However, I questioned the practical viability of the alternative mechanism he proposed.
This was a time when I had just learned about the school choice reforms in England and Chicago, and had begun systematically exploring potential applications of institution redesign with a minimalist mindset. My skepticism toward Switzer’s proposed mechanism stemmed from its departure from the principles of minimalist market design.
@@
From a purely academic standpoint, there is nothing wrong with this methodology. In fact, such policy recommendations are common under other institution design paradigms. However, in my view, this route posed a significant obstacle to the practical relevance of Switzer’s proposal.
Interpreting the proposed mechanism as interfering with an Army policy choice, I made the following point in my email reply:
“If I were to design a mechanism to correct the deficiency, I would have suggested exactly what you did in terms of the cadet preferences, but I would have kept the two types of positions separate. For the first type of positions, the priority would be based on cadet merit ranking, and for the second type of position, the priority would be adjusted based on whether the cadet is willing to pay the additional price or not. I think this proposal would have been ‘close’ to the current algorithm, and it would have still handled the deficiencies that you pointed out. In a way, I felt you adjusted the mechanism a little too much.”
Later, Switzer informed me that his proposal had indeed been dismissed by the relevant Army officers. The failures of the USMA-2006 mechanism identified in Switzer (2011), combined with my alternative proposal—the Multi-Price Cumulative Offer (MPCO) mechanism—led to a collaboration that resulted in Sönmez and Switzer (2013). [^91] MPCO mechanism is referred to as cadet-optimal stable mechanism (COSM) in Sönmez and Switzer (2013).

### v4

@@ 7.8.2 The Global Expansion of Kidney Exchange Since the Mid-2000s
After the first kidney exchange performed in the U.S. in Rhode Island in 2000, progress was initially slow. Between 2000 and 2003—prior to our engagement with the transplantation community—only 31 kidney exchange transplants were performed nationwide. In sharp contrast, a decade later the four-year total had risen to 2,078 between 2010 and 2013. The number climbed further over the next decade, with 4,387 kidney exchange transplants conducted between 2020 and 2023 (OPTN, 2025). A record 1,493 kidney exchange transplants were performed in 2024 alone, and by September 2025 the cumulative total in the U.S. exceeded 14,200 patients. (For annual transplant numbers in the U.S. from 2000 to 2024, see Figure 21.)
Figure 21: Annual number of kidney exchange transplants performed in the U.S. from 2000 to 2024, as reported in the UNOS database on August 27, 2025 (OPTN, 2025).
@@
According to estimates by Teltser (2019), 64% of these transplants—approximately 9,000—represent new transplants that would not have occurred without kidney exchange. Furthermore, Teltser (2019) highlights that kidney exchange has also improved average graft survival rates, despite the fact that kidney exchange recipients often come from more disadvantaged groups. These include highly sensitized patients with antibodies against large portions of the population, older patients, and recipients of kidneys from older donors.
@@
Over the past two decades, kidney exchange programs have expanded beyond the U.S., providing life-saving transplants to patients worldwide. Countries such as the UK, Turkey, Canada, India, Australia, Saudi Arabia, the Netherlands and Spain now perform some of the highest kidney exchange volumes outside the U.S. Altogether, the number of kidney exchange transplants carried out globally to date exceeds 23,000 (see Table 4 for worldwide kidney exchange activity since early 2000s).
Table 4: Expansion of kidney exchange activity worldwide since the early 2000s.
Country Program / Operator Start Documented volume
United States Three major platforms NKR, APKD, UNOS KPD & regional/hospital programs 2000 >>14,200   (by 08/2025)
Turkey Multiple single-center programs 2006 >>2,300   (by 12/2023)
United Kingdom UK Living Donor Kidney Sharing Scheme (NHS Blood & Transplant) 2007 >>2,000   (by 09/2024)
India Multiple single-center programs 2000 1,891   (09/2000–03/2024)
Canada Kidney Paired Donation Program 2009 >>1,000   (by 05/2023)
(Canadian Blood Services)
Australia Australian Paired Kidney Exchange Program 2010
New Zealand New Zealand Kidney Exchange 2011
Bilateral cooperation Australian and New Zealand Paired Kidney Exchange (ANZKX) 2019 >>600   (by 09/2025)
Saudi Arabia Kidney Paired Donation Program 2011 >>500   (by 10/2024)
(King Faisal Specialist Hospital )
Spain Organización Nacional de Trasplantes 2009 341   (by 12/2024)
Italy Centro Nazionale Trapianti 2006 91   (2005–2023)
Portugal Programa Nacional de Doação Renal Cruzada 2010 $\geq$24   (by 2020)
Trilateral cooperation Spain–Italy–Portugal cooperation protocol 2017 $\geq$19   (by 12/2024)
Netherlands National kidney exchange program 2004 284   (2004–2016)
Regional / single-center NDD chains 82   (by 04/2014)
Denmark
Finland Scandiatransplant Kidney Exchange 2019 95   (by Q2 2025)
Norway Programme (STEP)
Sweden
Austria 2015
Czech Republic Unified AT–CZ–IL program 2015 81   (2011–2019)
Israel 2019
Global Global Kidney Exchange (GKE) 2015 52   (2015–02/2022)
UK.
@@
The UK was one of the first countries to adopt market design tools for kidney exchange—including optimal matching algorithms—and now operates the largest national system. Since 2007, the National Health Service Blood and Transplant (NHSBT) has coordinated kidney exchanges through the UK Living Donor Kidney Sharing Scheme. This quarterly program has collaborated with market design researcher David Manlove since 2008 and celebrated its 2,000th kidney exchange transplant in September 2024.
@@
Initially limited to 2-way and 3-way cycles (Biró et al., 2009), the UK scheme has also been organizing non-directed donor (NDD) chains of up to three recipients since 2012 (Manlove and O’Malley, 2014; Ferrari et al., 2014). Among the 670 kidney exchange transplants conducted between April 2020 and March 2024, 126 (18.8%) were carried out in 2-way cycles, 118 (17.6%) in 3-way cycles, 215 (32.1%) in 2-way NDD chains, and 211 (31.5%) in 3-way NDD chains (NHS Blood and Transplant, 2024).
@@
Two key factors behind the UK’s success are the substantial participation of altruistic donors and compatible donor-recipient pairs. Between April 2015 and March 2024, there were 653 NDDs, with 347 donating directly to the deceased donor list and 306 donating through the kidney sharing scheme (NHS Blood and Transplant, 2024). Over this period, the proportion of NDDs participating in kidney exchanges grew from approximately 30% to 35% to 55% to 65%. Additionally, 119 compatible pairs participated in kidney exchanges during the same period.
@@
Taken together, the UK’s centrally managed national system—which incorporates up to 3-way cycles and NDD chains, employs an optimal matching algorithm, and benefits from significant participation by altruistic donors and compatible pairs—stands out as perhaps the closest large-scale realization of the vision pioneered in Roth et al. (2004, 2005b, 2006, 2007).
Turkey.
@@
Over 80% of kidney transplants in Turkey are from living donors, and the country led the world in living kidney donation rates in 2023, exceeding 35 donors per million people (IRODaT, 2024). Performing over 3,000 transplants annually, Turkey also ranks third globally in living-donor kidney transplants (LDKTs), behind only India and the U.S. These factors make Turkey an ideal candidate to implement a national system. Unfortunately, as discussed later in Section 12.4, my efforts with Ünver in this direction have not succeeded so far.
@@
Nevertheless, kidney exchange is widely practiced in Turkey through several single-hospital programs, with more than 2,300 transplants performed as of 2023. However, NDD chains are not utilized, as altruistic donation is prohibited by law.
@@
The Turkish Ministry of Health and the Turkish Society of Nephrology have tracked kidney exchange activity nationwide since 2012 (Turkish Society of Nephrology, 2023). According to their registries, 2,067 kidney exchange transplants were performed between 2012 and 2023, accounting for 6.3% of the country’s total LDKTs. [^98] Kidney exchange activity in Turkey predates 2012, with significant activity reported between 2006 and 2011. According to Yücetin et al. (2013), among the 59 hospitals performing LDKTs nationwide, Antalya Medicalpark Hospital, Istanbul Göztepe Medicalpark Hospital, and Akdeniz University Medical School Transplantation Center reported the highest kidney exchange volumes during this period. Together, they conducted 272 transplants, accounting for 14.4% of the LDKTs carried out at these centers. This relatively modest rate reflects the absence of a national kidney exchange system and the limited availability of kidney exchange in only some hospitals. However, in hospitals with kidney exchange programs, rates are significantly higher, around 15%. For example, Altun and Yavuz (2023) documented 141 kidney exchange transplants at Istanbul Göztepe Medicalpark Hospital from July 2011 to June 2021, representing 16.1% of the LDKTs performed at the center during that period.
India.
The most populous country in the world, India leads globally in LDKTs, performing over 10,000 in 2022 (IRODaT, 2024). Kidney exchange is widespread and conducted through numerous single-center programs across the country. Kute et al. (2025) reports that 1,891 kidney exchange transplants were performed between September 2000 and March 2024 across 84 transplant centers.
@@
As in Turkey, NDD chains are not practiced in India because altruistic donation is prohibited by law. Unlike in many other countries, kidney exchange operations in India are often carried out over several days, enabling the execution of large exchange cycles.
@@
At the single-center level, Kute et al. (2024) reports that the Dr HL Trivedi Institute of Transplantation Sciences in Ahmedabad—among the largest single-center kidney exchange programs in Asia—performed 539 kidney exchange transplants between January 2000 and March 2024. These included one 10-way and two 6-way exchanges conducted over multiple days, representing 10% of the 5,346 LDKTs performed at the center during this period. Similarly, Mahatma Gandhi Medical College in Jaipur, described by Mangal et al. (2024) as the largest program in Northern India, reported 127 kidney exchange transplants between 2014 and March 2024.
Canada.
The national Kidney Paired Donation program has been coordinated by Canadian Blood Services since 2009. The system runs every four months and incorporates cycles as well as NDD chains of up to four pairs. It is regarded as one of the most successful kidney exchange programs worldwide, largely because of the strong participation of altruistic donors. In May 2023, the program reached its 1,000th kidney exchange transplant, 68% of which were facilitated through donations from altruistic donors.
Saudi Arabia.
@@
More than 80% of kidney transplants in Saudi Arabia are from living donors, a pattern shared with Turkey and many other Asian countries. In 2023, the country ranked second worldwide—after Turkey—in the living kidney donation rate, with more than 35 donors per million people (IRODaT, 2024). The King Faisal Specialist Hospital and Research Centre celebrated its 500th kidney exchange transplant in October 2024 as part of its single-hospital Kidney Paired Donation Program, which has been in operation since 2011.
The Netherlands.
@@
National kidney exchange programs were launched in Australia in 2010 and in New Zealand in 2011, and the two were merged in 2019 into the Australian and New Zealand Paired Kidney Exchange (ANZKX) program. Since their inception, these programs have facilitated more than 600 kidney exchange transplants.
Spain / Italy / Portugal
@@
As another cross-border initiative, Spain, Italy, and Portugal signed a cooperation protocol in 2017 and achieved their first cross-border kidney exchange in 2018 (Valentín et al., 2019). Unlike other multi-country collaborations, however, each country continues to operate its own independent national system before engaging in cross-border exchanges.
The Organización Nacional de Trasplantes (ONT) launched Spain’s national kidney exchange program in 2009. Allowing 2-way and 3-way exchanges, it had facilitated 325 kidney exchange transplants by December 2023, with another 16 in 2024, bringing the total to 341 (Gobierno de España, 2024).
The Centro Nazionale Trapianti (CNT) introduced kidney exchange in Italy in 2006, making it one of the earliest adopters in Southern Europe. Between 2005 and 2023, 91 transplants were facilitated under the program, including 18 in 2023 alone (Centro Nazionale Trapianti, 2023).
The Programa Nacional de Doação Renal Cruzada (PNDRC) was established in Portugal in 2010. At least 24 transplants were reported through 2020 (Instituto Português do Sangue e da Transplantação, 2020), and more recent updates show continued activity, including several 2-way and 3-way cycles and two NDD-chain transplants in 2024.
By July 2023, a total of 16 patients had received transplants through the trilateral collaboration between Spain, Italy, and Portugal (Gobierno de España, 2024), with at least three additional international procedures recorded in 2024. This brought the cumulative cross-border total to more than 19 by the end of that year.
Scandinavia.
One of the most active cross-border collaborations in Europe is the Scandiatransplant Kidney Exchange Programme (STEP), with participation from Denmark, Finland, Norway, and Sweden (Weinreich et al., 2023). Launched in 2019 with the advocacy of economic designer Tommy Andersson and initially limited to 2-way exchanges, STEP has since been organizing 2-way and 3-way exchanges as well as NDD chains since January 2024. As of the second quarter of 2025, the program has facilitated a total of 95 transplants (Scandiatransplant, 2025).
Austria / Czech Republic / Israel.
Another cross-border kidney exchange partnership currently operates between Austria, the Czech Republic, and Israel. Due to the limited size of their national pools, Austria and the Czech Republic—both of which introduced national kidney exchange systems in 2011—merged their programs in 2015. This collaboration led to the first cross-border kidney exchange in 2016 (Böhmig et al., 2017) and was further expanded to include Israeli centers in 2019. Between 2011 and 2019, these systems facilitated 81 transplants, including 33 after 2015 under the unified program, of which eight were via cross-border exchanges (Viklicky et al., 2020).
Global Kidney Exchange.
Finally, a related cross-border initiative is the Global Kidney Exchange (GKE) program, spearheaded by Alvin Roth and Michael Rees (Rees et al., 2017, 2022). The program funds kidney transplants for recipients from low- and middle-income countries in exchange for living donors who facilitate transplant cycles or NDD chains in high-income countries. Since 2015, the Alliance for Paired Kidney Donation (APKD) has implemented GKE. Between January 2015 and February 2022, 17 international patients facing financial barriers to direct transplantation facilitated kidney transplants for 35 U.S. patients through GKE, resulting in a total of 52 kidney transplants. [^99] Our earlier efforts on kidney exchange in New England and elsewhere aligned with minimalist principles and, most importantly, strictly adhered to the ethical guidelines of the transplantation community as codified in the Consensus Statement (Abecassis et al., 2000). Consequently, these initiatives did not encounter major opposition within the transplantation community. In contrast, the GKE program has sparked debate, receiving both endorsements and criticisms. Supporters include prominent bioethicists such as Peter Singer (Minerva et al., 2019) and the Italian government, which has recommended GKE to the World Health Organization (Pullen, 2018). However, unlike other kidney exchange initiatives, GKE has also faced strong ethical objections. Opponents include Frank Delmonico, who led criticism of GKE in his role as Executive Director of the Declaration of Istanbul Custodian Group, as well as the European Union National Competent Authorities, which oversee transplant activities across the EU (Delmonico and Ascher, 2017; DOICG, 2017; EU-NCA, 2018). For an in-depth discussion of this debate, see Ambagtsheer et al. (2020).
European Union.
@@
ENCKEP led to a follow-on project called Software for Transnational Kidney Exchange Programmes (KEP-SOFT), funded by a COST Innovators Grant from November 2021 to October 2022. This effort produced KEPsoft, a software solution for national and international kidney exchange (KEP-SOFT, 2024).100100100KEPsoft was jointly developed by the University of Glasgow, INESC TEC (Porto), HUN-REN-KRTK (Budapest), and Óbuda University (Budapest) (KEP-SOFT, 2024).
Further affirming the growing enthusiasm for kidney exchange in Europe, the European Commission launched a 2023 call for proposals for action grants under its Health Systems and Healthcare Workforce initiative, specifically targeting “Facilitating Organ Paired Exchange” (European Commission, 2023). This initiative aims to harmonize organ donation and transplantation practices across EU member states by developing a shared algorithm for matching donor-patient pairs and establishing common protocols, platforms, and governance frameworks. The ultimate goal is to create an EU-wide system that expands kidney exchange opportunities and improves patient outcomes.101101101The EURO-KEP project, led by the Spanish National Transplant Organization (ONT), was the sole project funded under this call and is supported by KEPsoft software (KEP-SOFT, 2024).
@@
My holistic approach to research and policy in school choice (see Section 6) and kidney exchange (see Section 7) was primarily driven by instinct. Early in my academic career, I did not consciously follow any specific institutional design framework. In both applications, my goal was to address the limitations of existing resource allocation systems in ways that could gain acceptance from decision-makers vested in preserving them.
Clarity emerged as I observed parallels between the research-guided school choice reform in Boston (see Section 6.10) and the policy-maker-driven reforms in England and Chicago (see Section 6.12). It became evident that the methodology evolving from my holistic research and policy efforts could provide the foundation for a promising economic design framework on a broader scale.
I interpreted these developments in England and Chicago as providing external validity for this emerging methodology. My approach—uncovering how real-life institutions had diverged from their intended roles and restoring their alignment—resonated with decision-makers. This emphasis on aligning institutions with their original intentions enabled me to achieve policy impact, which may explain my success.
@@
In May 2011, I received an unexpected email from (then) Major Tobias Switzer, a U.S. Air Force officer, seeking feedback on his master’s thesis (Switzer, 2011) from the Catholic University of Chile. Following the basic structure of Balinski and Sönmez (1999) and adopting the no justified envy axiom (see Definition 5) for his framework, Switzer examined the branching process of cadets into military specialties at the United States Military Academy (USMA).102102102Following the terminology introduced in Balinski and Sönmez (1999), Switzer (2011) also refers to the no justified envy axiom as fairness.
Upon reading his thesis, I found the problem fascinating and his critique of the mechanism used at West Point—henceforth the USMA-2006 mechanism—to be valid. Little did I know that, building on the use-inspired theory in Balinski and Sönmez (1999), his thesis would mark a key step in yet another “discovery–invention cycle” (see Figure 36).
Despite my enthusiasm, however, I questioned the practical viability of his proposed alternative mechanism. At the time, I had just learned about the school choice reforms in England and Chicago and had begun systematically exploring potential applications of institutional redesign with a minimalist mindset. My skepticism toward Switzer’s proposal stemmed from its departure from the principles of minimalist market design.
@@
From a purely academic standpoint, there is nothing wrong with this methodology. In fact, such policy recommendations are common under other economic design paradigms. However, in my view, this route posed a significant obstacle to the practical relevance of Switzer’s proposal.
Interpreting the proposed mechanism as interfering with an Army policy choice unrelated to the underlying shortcomings Switzer identified, I made the following point in my email reply:
“If I were to design a mechanism to correct the deficiency, I would have suggested exactly what you did in terms of the cadet preferences, but I would have kept the two types of positions separate. For the first type of positions, the priority would be based on cadet merit ranking, and for the second type of position, the priority would be adjusted based on whether the cadet is willing to pay the additional price or not. I think this proposal would have been close’ to the current algorithm, and it would have still handled the deficiencies that you pointed out. In a way, I felt you adjusted the mechanism a little too much.”
Later, Switzer informed me that his proposal had indeed been dismissed by the relevant Army officers. The failures of the USMA-2006 mechanism identified in Switzer (2011), combined with my alternative proposal—the Multi-Price Cumulative Offer (MPCO) mechanism—led to a collaboration that resulted in Sönmez and Switzer (2013).103103103MPCO mechanism is referred to as cadet-optimal stable mechanism (COSM) in Sönmez and Switzer (2013).